We want to determine the AOQ for an acceptance sampling plan when the quality of the incoming lots in percent defective is 4​%, and then again when the incoming percent defective is 8.5​%. The sample size is 50 units for a lot size of 600 units. Furthermore, Pa at 4​% defective levels is 0.89. At 8.5​% incoming defective​ levels, the Pa is found to be 0.59. Determine the average outgoing quality for both incoming percent defective levels.

The average outgoing quality for 8.5​% incoming percent defective level is __. ​(Round your response to three decimal​ places.)
Answer #1

The average outgoing quality for 4​% incoming percent defective level is 0.033

The average outgoing quality for 8.5​% incoming percent defective level is 0.046

Explanation:

Average outgoing quality or AOQ for percent defective= (Pd)(Pa)(N-n)/N

Pd= true percent defective of the lot = 4%= 0.04

Pa = probability of accepting the lot = 0.89

N= number of items in the lot = 600

n= number of items in the sample= 50

Hence AOQ = 0.04*0.89*(600-50)/600 = 0.033

When Pd = 8.5%= 0.085 and Pa = 0.59,

AOQ = 0.085*0.59*(600-50)/600 = 0.046